# KKL theorem for the influence of a set of variables

### v1

@@ Abstract.
We study the notion of the influence of a set of variables on a Boolean function, which was recently introduced by Tal. We show that for an arbitrary fixed $d$, every Boolean function $f$ on $n$ variables admits a $d$-set of influence at least $\tfrac{1}{10}\mathbf{W}^{\geqslant d}(f)(\frac{\log n}{n})^{d}$, which is a direct generalisation of the Kahn–Kalai–Linial theorem. We give an example demonstrating essential sharpness of this result. Further, we generalise a related theorem of Oleszkiewicz regarding influences of pairs of variables.
@@ 1. Introduction
The analysis of Boolean functions is a vibrant area of research at the intersection of discrete mathematics and computer science. In this work, we study a multi-bit generalization of a central notion in the field—influence.
Let us lay the foundation of our considerations. For $\mathbf{i}\subseteq[n]:=\{1,2,\ldots,n\}$ define the Walsh functions $\chi_{\mathbf{i}}\colon\{-1,1\}^{n}\to\mathbb{R}$ by $\chi_{\mathbf{i}}(x):=\prod_{j\in\mathbf{i}}x_{j}$ and $\chi_{\varnothing}(x):=1$. The system $\{\chi_{\mathbf{i}}\colon\mathbf{i}\subseteq[n]\}$ is an orthonormal basis of $L^{2}(\{-1,1\}^{n},\mu_{n})$, which yields that every function $f\colon\{-1,1\}^{n}\to\mathbb{R}$ has a unique Fourier–Walsh expansion $f=\sum_{\mathbf{i}\subseteq[n]}\hat{f}(\mathbf{i})\chi_{\mathbf{i}}$, where $\hat{f}(\mathbf{i})\in\mathbb{R}$.
For $i\in[n]$ define the (single-bit) influence of bit $i$ on $f$ by $\operatorname{Inf}_{i}(f):=\sum_{\mathbf{j}\subseteq[n]:\ \!i\in\mathbf{j}}\hat{f}(\mathbf{j})^{2}.$ This is a well-studied notion; its history and various properties can be found, alongside many other topics this paper is touching upon, in the textbook of Ryan O’Donnel [9]. A celebrated result regarding the influence is the Kahn–Kalai–Linial theorem [6], which says that every Boolean function admits a very influential bit. Let $\mathbf{W}^{\geqslant d}(f):=\sum_{\mathbf{i}\subseteq[n]\colon\!|\mathbf{i}|\geqslant d}\hat{f}(\mathbf{i})^{2}$.
@@
$\operatorname{Inf}_{i}(f)\geqslant C_{1}\mathbf{W}^{\geqslant 1}(f)\frac{\ln n}{n}.$
@@
Several generalizations of the definition of the single-bit influence accounting for multiple bits have been considered, see [2]. One such generalization has been recently introduced by Tal [11] and is (equivalent to) the following.
Definition 1.2.
Let $f\colon\{-1,1\}^{n}\to\mathbb{R}$ and $\mathbf{i}\subseteq[n]$. We define the influence of $\mathbf{i}$ on $f$ by
@@
Approximately speaking, $\operatorname{Inf}_{\mathbf{i}}(f)$ measures if all the bits of $\mathbf{i}$ are influential on $f$. This stays in contrast with other multi-bit generalization of influence, studied by Ben-Or and Linial [1], which measures if any bit of $\mathbf{i}$ is influential. For a more precise discussion, see Section 3.2.
Every definition is as good as how useful it is or how much one can prove about it, and therefore let us discuss some relevant results. A close variant of Definition 1.2 in case of $|\mathbf{i}|=2$ has been independently considered by Tanguy [12]. He obtained a result in the spirit of the KKL theorem. Namely, he proved that every Boolean $f$ either admits a pair of bits with influence at least $C\mathbf{W}^{\geqslant 1}(f)(\frac{\ln n}{n})^{2}$, or there is a single bit with influence at least $C(\mathbf{W}^{\geqslant 1}(f)/n)^{\theta}$, for some universal constants $C>0$ and $\theta\in(1/2,1)$.
@@
where $C_{KS}$ is a constant depending only on $d$.
In the course of the proof of Theorem 1.3, Oleszkiewicz shows that every Boolean function $f$ admits a pair of distinct bits $i$$j$ such that $\operatorname{Inf}_{\{i,j\}}(f)\geqslant C_{O}\mathbf{W}^{\geqslant 2}(f)(\frac{\ln n}{n})^{2}$. Hence the assumption in Theorem 1.3 of small pair influences yields that $\mathbf{W}^{\geqslant 2}(f)$ is small, so in particular implies the assumption of Theorem 1.5 for $d=1$. However, the stronger assumption in Theorem 1.3 yields the stronger consequence of extra $\ln n/n$ factor.
@@ 1.1. Main results
$\operatorname{Inf}_{\mathbf{i}}(f)\geqslant\frac{1}{10}\mathbf{W}^{\geqslant d}(f)\left(\frac{\ln n}{n}\right)^{d},$
@@
Observe that for $d=1$ we recover the KKL theorem, and the case $d=2$ is the above-mentioned consequence of the proof of Theorem 1.3. We prove Theorem 1.6 in Section 4 by generalising the representation of $\mathbf{W}^{\geqslant 2}(f)$ and the hypercontractivity lemma from [10].
In Section 5 (Theorem 5.10), we present a family of functions showing that for every fixed $d$ Theorem 1.6 is asymptotically sharp (up to the $1/10$ constant). The constructed examples, which we call the $d$-hypertribes, are a variation on the classical tribe functions due to Ben-Or and Linial [1] which demonstrate the sharpness of the KKL theorem. As was pointed out in [10], the standard tribe function cannot serve as a sharp example. E.g. for $d=2$ influences of pairs of bits from the same tribe are larger than the influence of a typical pair. To account for this, one needs to increase the size of tribes and allow voters to belong to multiple tribes in a controlled way. These adjustments complicate the structure of the function, and the explanation why it serves as a sharp example becomes far from trivial. The construction of the $d$-hypertribes was suggested to us by Peter Keevash in a private communication.
@@
for some $\alpha\in(0,C_{3})$ and all $\mathbf{i}\subseteq[n]$ of size $d+1$, then there exists $g\colon\{-1,1\}^{n}\to\{-1,1\}$ of degree $d$ such that
@@
For $d=1$ we recover Theorem 1.3. Indeed, the only $1$-degree Boolean functions are the constants $-1$, $+1$, single-bit functions $\chi_{\{i\}}\equiv x_{i}$ and their negations. Depending which of these $g$ is, from Theorem 1.7 we can recover the relevant alternative of Theorem 1.3.
We will see in Section 3 that $\operatorname{Inf}_{i}(f)=\|\partial_{\mathbf{i}}f\|_{2}^{2}$, where $\partial_{\mathbf{i}}f$ is a (discrete) $d$-order partial derivative of $f$. From this point of view, Theorem 1.7 is a Boolean variant of a theorem from real analysis stating roughly that if all the $(d+1)$-order partial derivatives of $f\colon[0,1]^{n}\to\mathbb{R}$ are small, then $f$ is close to some polynomial of degree at most $d$.
@@
Let $n>d\geqslant 1$. Assume that $f\colon\{-1,1\}^{n}\to\{-1,1\}$ satisfies (1) and that for some $l\in[d]$ we have $|\hat{f}(\mathbf{i})|\leqslant 2^{-d}$ for all $\mathbf{i}\subseteq[n]$ of size at most $l-1$. Then there exists a function $g\colon\{-1,1\}^{n}\to\{-1,1\}$ of $d$-degree such that
@@ 2. Notations
Throughout the paper, $n$ is an arbitrary positive integer (possibly further constrained). Let $d\in\mathbb{Z}_{\geqslant 0}$. We use the standard notation $[n]=\{1,2,\ldots,n\}$ and the less standard notation
$[n]_{d}:=\{A\subseteq[n]\colon|A|=d\},\quad[n]_{<d}:=\{A\subseteq[n]\colon|A|<d\}.$
We define $[n]_{\leqslant d}$, $[n]_{>d}$ and $[n]_{\geqslant d}$ in an analogous manner.
We write $\operatorname{Im}(f)$ for the image of a function $f$, and refer to functions with $\operatorname{Im}(f)\subseteq\{-1,1\}$ as Boolean functions. Plain letters $i$, $j$, $k$ usually denote indices of bits of the input, i.e. elements of $[n]$, whereas bold letters $\mathbf{i}$, $\mathbf{j}$, $\mathbf{k}$ refer to index sets, i.e. subsets of $[n]$. Sometimes, for clarity, we omit brackets and write e.g. $\mathbf{i}\setminus i$ instead of $\mathbf{i}\setminus\{i\}$. The name $d$-set will stand for a $d$ element subset of $[n]$.
@@ 3. Basics of harmonic analysis
The probabilities $\operatorname{\mathbb{P}}$ and expectations $\operatorname{\mathbb{E}}$ are always taken with respect to the measure $\mu_{n}$. The bold letter $\mathbf{x}$ will then denote a uniform random variable with distribution $\mu_{n}$. Note that, due to the product structure of $\mu_{n}$, the set of single bits $\mathbf{x}_{i}$ of $\mathbf{x}$ is a family of independent and identically distributed random variables, each having distribution $\frac{1}{2}\delta_{-1}+\frac{1}{2}\delta_{1}$. Often, when it is clear from the context, we omit the letter $\mathbf{x}$ in probabilities or expectations.
@@
Further, note that we have $\operatorname{\mathbb{E}}f=\langle f,1\rangle=\langle f,\chi_{\varnothing}\rangle=\hat{f}(\varnothing)$. Moreover, if $f$ is Boolean (so $\operatorname{Im}{f}\subseteq\{-1,1\}$) we have $\|f\|_{2}^{2}=\sum_{\mathbf{i}\subseteq[n]}\hat{f}(\mathbf{i})^{2}=1$ and so in this case $|\hat{f}(\mathbf{i})|\leqslant 1$ for every $\mathbf{i}\subseteq[n]$.
@@ 3.1. Derivative
$\partial_{i}f(x):=\frac{f(x^{i\to+1})-f(x^{i\to-1})}{2}$
where for $\varepsilon\in\{-1,1\}$,
$x^{i\to\varepsilon}:=(x_{1},\ldots,x_{i-1},\varepsilon,x_{i+1},\ldots,x_{n}).$
@@
$\partial_{\mathbf{i}}f(x)=\frac{1}{2^{|\mathbf{i}|}}\sum_{\varepsilon\in\{-1,1\}^{\mathbf{i}}}\Big{(}f(x^{\mathbf{i}\to\varepsilon})\cdot\prod_{j\in\mathbf{i}}\varepsilon_{j}\Big{)},$ (3)
where $\varepsilon_{j}$ are the bits of $\varepsilon$ and $x^{\mathbf{i}\to\varepsilon}\in\{-1,1\}^{n}$ is defined via
$(x^{\mathbf{i}\to\varepsilon})(j):=\begin{cases}\varepsilon_{j},&\text{if}\ j\in\mathbf{i},\\ x_{j},&\text{if}\ j\notin\mathbf{i}.\end{cases}$
@@
Proposition 3.4.
Let $f\colon\{-1,1\}^{n}\to\mathbb{R}$ and $\mathbf{i}\subseteq[n]$ satisfy $\partial_{\mathbf{i}}f(x)\neq 0$ for some $x\in\{-1,1\}^{n}$. Then for every $i\in\mathbf{i}$ there exists $\varepsilon\in\{-1,1\}^{\mathbf{i}\setminus i}$ such that $f(x^{\mathbf{i}\to\varepsilon^{+}})\neq f(x^{\mathbf{i}\to\varepsilon^{-}})$, where $\varepsilon^{\pm}\in\{-1,1\}^{\mathbf{i}}$ is defined by
$\varepsilon^{\pm}_{j}:=\begin{cases}\varepsilon_{j},\ &\text{if}\ j\neq i,\\ \pm 1,\ &\text{if}\ j=i.\end{cases}$
Proof.
By Proposition 3.2 we have
$=\frac{1}{2^{|\mathbf{i}|}}\sum_{\varepsilon\in\{-1,1\}^{\mathbf{i}\setminus i}}\Big{(}f(x^{\mathbf{i}\to\varepsilon^{+}})-f(x^{\mathbf{i}\to\varepsilon^{-}})\Big{)}\prod_{j\in\mathbf{i}\setminus i}\varepsilon_{j}.$
Since $\partial_{\mathbf{i}}f(x)\neq 0$, at least one of the summands must be non-zero, and the statement follows. ∎

### v2

@@ Abstract.
Consider a Boolean function $f$ on the $n$-dimensional hypercube, and a set of variables (indexed by) $S\subseteq\{1,2,\ldots,n\}.$ The coalition influence of the variables $S$ on a function $f$ is the probability that after a random assignment of variables not in $S$, the value of $f$ is undetermined. In this paper, we study a complementary notion, which we call the joint influence: the probability that, after a random assignment of variables not in $S$, the value of $f$ is dependent on all variables in $S$.
We show that for an arbitrary fixed $d$, every Boolean function $f$ on $n$ variables admits a $d$-set of joint influence at least $\tfrac{1}{10}\mathbf{W}^{\geqslant d}(f)(\frac{\log n}{n})^{d}$, where $\mathbf{W}^{\geqslant d}(f)$ is the Fourier weight of $f$ at degrees at least $d$. This result is a direct generalisation of the Kahn–Kalai–Linial theorem. Further, we give an example demonstrating essential sharpness of the above bound. In our study of the joint influence we consider another notion of multi-bit influence recently introduced by Tal.
@@ 1. Introduction
Consider $i\in[n]=\{1,2,\ldots,n\}$ and $f\colon\{-1,1\}^{n}\to\{-1,1\}$. Throughout, we consider the uniform measure $\mu_{n}$ on $\{-1,1\}^{n}$. Define the influence of bit $i$ on $f$ by
$\operatorname{Inf}_{i}(f):=\underset{\mathbf{x}\sim\mu_{n}}{\operatorname{\mathbb{P}}}(f(\mathbf{x})\neq f(\mathbf{x}^{\oplus i})),$
where $x^{\oplus i}=(x_{1},\ldots,x_{i-1},-x_{i},x_{i+1},\ldots,x_{n})$.
The influence quantifies how often a given bit can sway the outcome of the function. It is a well-studied notion central to the field of Boolean analysis. Its history and various properties can be found in the textbook of Ryan O’Donnell [12]. A celebrated result, due to Kahn, Kalai and Linial [8], says that every Boolean function admits a very influential bit.
@@
$\operatorname{Inf}_{i}(f)\geqslant C_{1}\operatorname{\mathbb{V}\!\mathrm{a}\mathrm{r}}(f)\frac{\ln n}{n},$
where $\operatorname{\mathbb{V}\!\mathrm{a}\mathrm{r}}(f)=\underset{\mathbf{x}\sim\mu_{n}}{\operatorname{\mathbb{V}\!\mathrm{a}\mathrm{r}}}(f(\mathbf{x})).$
@@
Several generalizations of the definition of the single-bit influence accounting for multiple bits have been considered, see [3] for discussion. The one probably most studied is the coalition influence, which we will now define.
In what follows, for brevity, we associate a variable $x_{i}\in\{-1,1\}$ with its index $i$. Let $\mathbf{i}\subseteq[n]$ be a set of variables. The coalition influence $\operatorname{CInf}_{\mathbf{i}}(f)$ of $\mathbf{i}$ on $f$ is defined as the probability that assigning values to the variables not in $\mathbf{i}$ at random, the value of $f$ is undetermined. There has been a long line of research studying this type of influence, for the main results see [8], [1], [5].
The focus of our paper is a counterpart notion to the coalition influence. We call it the joint influence $\operatorname{JInf}_{\mathbf{i}}(f)$ of $\mathbf{i}$ on $f$ and define it as the probability that assigning values to the variables not in $\mathbf{i}$ at random, the value of $f$ is dependent on all of the variables in $\mathbf{i}$.
More formally, say that $i\in\mathbf{i}$ is $\mathbf{i}$-pivotal for $f$ on input $x$ if there is an input $y$ agreeing with $x$ on bits $n]\setminus\mathbf{i}$ such that $f(y)\neq f(y^{\oplus i})$. Then for non-empty $\mathbf{i}$ define
[…]
Note that using $\mathbf{i}$-pivotality, the coalition influence satisfies
[…]
In particular, $\operatorname{CInf}$ can be thought of as a union of influences, whereas $\operatorname{JInf}$ as an intersection, and hence the name. As a result, it follows that $\operatorname{CInf}_{\mathbf{i}}(f)\geqslant\operatorname{JInf}_{\mathbf{i}}(f)$.
Let us mention yet another generalization of the notion of single bit influence, purely analytic one, as it will turn out useful in studying the joint influence. For $\mathbf{i}\subseteq[n]$ we define the Walsh functions $\chi_{\mathbf{i}}\colon\{-1,1\}^{n}\to\mathbb{R}$ by $\chi_{\mathbf{i}}(x):=\prod_{j\in\mathbf{i}}x_{j}$ and $\chi_{\varnothing}(x):=1$. The system $\{\chi_{\mathbf{i}}\colon\mathbf{i}\subseteq[n]\}$ is an orthonormal basis of $L^{2}(\{-1,1\}^{n},\mu_{n})$, which yields that every function $f\colon\{-1,1\}^{n}\to\mathbb{R}$ has a unique Fourier–Walsh expansion $f=\sum_{\mathbf{i}\subseteq[n]}\hat{f}(\mathbf{i})\chi_{\mathbf{i}}$, where $\hat{f}(\mathbf{i})\in\mathbb{R}$. We define the T-influence $\operatorname{Inf}_{\mathbf{i}}(f)$ by
@@
It has been introduced by Tal in [14] in his study of the Fourier mass at high levels in $\mathbb{AC^{0}}$ Boolean circuits. However, it has been independently considered by Tanguy in [15] for two-element $\mathbf{i}$’s and in this case further studied by Oleszkiewicz in [13]. Often, when clear from the context, we will refer to it simply as the influence. Sometimes, we will say that $\operatorname{Inf}_{\mathbf{i}}(f)$ is a $d$-degree influence, where $d=|\mathbf{i}|$.
The T-influence, although interesting on its own, will open to us the Fourier methods to bound the joint influence via inequality $\operatorname{JInf}_{\mathbf{i}}(f)\geqslant\operatorname{Inf}_{\mathbf{i}}(f)$, proven in Proposition 3.5. Further, note that in the case of $\mathbf{i}=\{i\}$ all the three generalization of the single-bit influence are equal to $\operatorname{Inf}_{i}(f)$ (with the equality $\operatorname{Inf}_{\{i\}}(f)=\operatorname{Inf}_{i}(f)$ being folklore).
1.1. T-influence results
Let us say a few words on what relevant to us has been proved on the T-influence.
The case of $|\mathbf{i}|=2$ has been considered by Tanguy in [15], where he obtained a result in the spirit of the KKL theorem. Namely, he proved that every Boolean $f$ either admits a pair of bits with T-influence at least $C\operatorname{\mathbb{V}\!\mathrm{a}\mathrm{r}}(f)(\frac{\ln n}{n})^{2}$, or there is a single bit with influence at least $C(\operatorname{\mathbb{V}\!\mathrm{a}\mathrm{r}}(f)/n)^{\theta}$, for some universal constants $C>0$ and $\theta\in(1/2,1)$.
@@ 1.1. T-influence results
where $C_{KS}$ is a constant depending only on $d$ and $\mathbf{W}^{\geqslant d+1}(f)=\sum_{\mathbf{i}\subseteq[n]\colon\!|\mathbf{i}|\geqslant d+1}\hat{f}(\mathbf{i})^{2}$.
In the course of the proof of Theorem 1.2, Oleszkiewicz shows that every Boolean function $f$ admits a pair of distinct bits iijj such that $\operatorname{Inf}_{\{i,j\}}(f)\geqslant C_{O}\mathbf{W}^{\geqslant 2}(f)(\frac{\ln n}{n})^{2}$. Hence the assumption in Theorem 1.2 of small pair influences yields that $\mathbf{W}^{\geqslant 2}(f)$ is small, so in particular implies the assumption of Theorem 1.4 for $d=1$. However, the stronger assumption in Theorem 1.2 yields the stronger consequence of extra $\ln n/n$ factor.
@@ 1.2. Main results
$\operatorname{JInf}_{\mathbf{i}}(f)\geqslant\operatorname{Inf}_{\mathbf{i}}(f)\geqslant\frac{1}{10}\mathbf{W}^{\geqslant d}(f)\left(\frac{\ln n}{n}\right)^{d},$
@@
Observe that for $d=1$ we recover the KKL theorem, and the case $d=2$ is the above-mentioned consequence of the proof of Theorem 1.2. We prove Theorem 1.5 in Section 4 by generalising the representation of $\mathbf{W}^{\geqslant 2}(f)$ and the hypercontractivity lemma from [13]. Recall that the first inequality holds actually for every $\mathbf{i}$ thanks to Proposition 3.5.
In Section 5 we prove that Theorem 1.5 is asymptotically sharp in the following sense.
Theorem 1.6.
For every $d\geqslant 2$ there is a constant $C^{d}_{H}$ depending only on $d$, and a family of functions $H_{n}^{d}$ such that for every $\mathbf{i}\subseteq[n]$ of size $d$ we have
$\operatorname{Inf}_{\mathbf{i}}(H_{n}^{d})\leqslant\operatorname{JInf}_{\mathbf{i}}(H_{n}^{d})\leqslant C^{d}_{H}\mathbf{W}^{\geqslant d}(H_{n}^{d})\left(\frac{\log_{2}n}{n}\right)^{d}.$
The functions $H_{n}^{d}$, which we call the $d$-hypertribes, are a variation on the classical tribe functions due to Ben-Or and Linial [2] demonstrating the sharpness of the KKL theorem. As was pointed out in [13], the standard tribe function cannot serve as a sharp example in our case. E.g. for $d=2$ influences of pairs of bits from the same tribe are larger than the influence of a typical pair. To account for this, one needs to increase the size of tribes and allow voters to belong to multiple tribes in a controlled way. These adjustments complicate the structure of the function, and the explanation why it serves as a sharp example becomes far from trivial. The construction of the $d$-hypertribes was suggested to us by Peter Keevash in a private communication.
@@
for all $\mathbf{i}\subseteq[n]$ of size $d+1$ for some $\alpha\in(0,C_{3})$, then there exists $g\colon\{-1,1\}^{n}\to\{-1,1\}$ of degree $d$ such that
@@
For $d=1$ we recover Theorem 1.2. Indeed, the only $1$-degree Boolean functions are the constants 1-1, +1+1, single-bit functions $\chi_{\{i\}}\equiv x_{i}$ and their negations. Depending which of these $g$ is, from Theorem 1.7 we can recover the relevant alternative of Theorem 1.2.
We will see in Section 3 that $\operatorname{Inf}_{\mathbf{i}}(f)=\|\partial_{\mathbf{i}}f\|_{2}^{2}$, where $\partial_{\mathbf{i}}f$ is a (discrete) $d$-order partial derivative of $f$. From this point of view, Theorem 1.7 is a Boolean variant of a theorem from real analysis stating roughly that if all the $(d+1)$-order partial derivatives of $f\colon[0,1]^{n}\to\mathbb{R}$ are small, then $f$ is close to some polynomial of degree at most $d$.
@@
Let $n>d\geqslant 1$. Assume that $f\colon\{-1,1\}^{n}\to\{-1,1\}$ satisfies (1) and that for some $l\in[d]$ we have $|\hat{f}(\mathbf{i})|\leqslant 2^{-d}$ for all $\mathbf{i}\subseteq[n]$ of size at most l1l-1. Then there exists a function $g\colon\{-1,1\}^{n}\to\{-1,1\}$ of $d$-degree such that
@@ 2. Notations
Throughout the paper, $n$ is an arbitrary positive integer (possibly further constrained). Let $d\in\mathbb{Z}_{\geqslant 0}$. We use the standard notation $n]=\{1,2,\ldots,n\}$ and the less standard notation
$n]_{d}:=\{A\subseteq[n]\colon|A|=d\},\quad[n]_{<d}:=\{A\subseteq[n]\colon|A|<d\}.$
We define $n]_{\leqslant d}$, $n]_{>d}$ and $n]_{\geqslant d}$ in an analogous manner.
We write $\operatorname{Im}(f)$ for the image of a function $f$, and refer to functions with $\operatorname{Im}(f)\subseteq\{-1,1\}$ as Boolean functions. Plain letters $i$, $j$, $k$ usually denote indices of bits of the input, i.e. elements of $[n]$, whereas bold letters $\mathbf{i}$, $\mathbf{j}$, $\mathbf{k}$ refer to index sets, i.e. subsets of $[n]$. For brevity, we will associate a variable $x_{i}\in\{-1,1\}$ with its index $i$, and so we will say e.g. variables $\mathbf{i}$ instead of variables $\{x_{i}\colon i\in\mathbf{i}\}$. Sometimes, for clarity, we omit brackets and write e.g. $\mathbf{i}\setminus i$ instead of $\mathbf{i}\setminus\{i\}$. The name $d$-set will stand for a $d$ element subset of $[n]$.
For an input $x=(x_{1},\ldots,x_{n})$, $i\in[n]$ and $\varepsilon\in\{-1,1\}$ we write
$x^{\oplus i}$ $=(x_{1},\ldots,x_{i-1},-x_{i},x_{i+1},\ldots,x_{n}),$
$x^{i\to\varepsilon}$ $=(x_{1},\ldots,x_{i-1},\varepsilon,x_{i+1},\ldots,x_{n}),$
and further, for $\mathbf{i}\subseteq[n]$ and $y\in\{-1,1\}^{\mathbf{i}}$ define $x^{\mathbf{i}\to y}\in\{-1,1\}^{n}$ by
$x^{\mathbf{i}\to y}_{j}=\begin{cases}x_{j},\text{ if }j\in[n]\setminus\mathbf{i},\\ y_{j},\text{ if }j\in\mathbf{i}.\end{cases}$
@@ 3. Basics of harmonic analysis
[…]
@@
[…]
@@ 3.1. Derivative
$\partial_{i}f(x):=\frac{f(x^{i\to+1})-f(x^{i\to-1})}{2}.$
@@
$\partial_{\mathbf{i}}f(x)=\frac{1}{2^{|\mathbf{i}|}}\sum_{y\in\{-1,1\}^{\mathbf{i}}}\Big{(}f(x^{\mathbf{i}\to y})\cdot\prod_{j\in\mathbf{i}}y_{j}\Big{)}.$ (3)